OpenStudy (anonymous):

Find the altitude and the area of an equilateral triangle whose side is 8 cm long.

OpenStudy (gorv):

in equilateral traingle altitude will be perpendicular to base and bisect it

OpenStudy (gorv):



OpenStudy (gorv):

from pythagoras theorm altitude=\[\sqrt{8^{2}-4^{2}}=\sqrt{64-16}=\sqrt{48}=\sqrt{16*3}=4*\sqrt{3}\]

OpenStudy (gorv):

area=(altitude*base)/2 \[\frac{ 4*\sqrt{3}*8 }{ 2 }=16*\sqrt{3}\]

OpenStudy (gorv):

@conan.195114

OpenStudy (anonymous):

sorry, used\[h=\frac{ 2\sqrt{s(s-a)(s-b)(s-c)} }{c}\] and got \[4\sqrt{3}\] is this correct?

OpenStudy (anonymous):

for area i used the heron's formula and got the same answer \[16\sqrt{3}\] thanks for helping me though i found it too long to wait so i studied on my own. thanks again

OpenStudy (gorv):

okk u can also do so ....but in heron calculation will be more